OpenStudy (anonymous):

please help with this... Consider the function z(x, y) = ye^(2x^3−5) at (x, y) = (1, 3). Use the property of partial derivative to find how the number of units x must approximately change, if y increases by 0.5 unit so that z increases by 1 unit. i got this till now Z'x=y.6.e^2x^3-5 z'y= y.e^(2x^3-5)

OpenStudy (anonymous):

\[z'_{y} = e ^{2x^3-5}\] my bad

ganeshie8 (ganeshie8):

\[\large \mathbb{ \Delta x = x_y\Delta y+x_z \Delta z}\]

ganeshie8 (ganeshie8):

thats the approximation formula right ?

OpenStudy (anonymous):

i have no idea tbh =/

ganeshie8 (ganeshie8):

have u seen total differential formula : \[\large df = \dfrac{\partial f}{\partial x}dx + \dfrac{\partial f}{\partial y}dy \] ?

OpenStudy (anonymous):

that's just the total differentiation right??

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

oh ok, so u just rewrote the total differentation to get dx ?

ganeshie8 (ganeshie8):

exactly! and i have replaced differentials with Delta

OpenStudy (anonymous):

thnx! gonna try rewrite it myself to see if i get there, btw are my derivatives correct?

ganeshie8 (ganeshie8):

\[\large z = ye^{2x^3-5} \] is this the function ?

OpenStudy (anonymous):

yup

ganeshie8 (ganeshie8):

when u differentiate with respect to x, all other independent variables can be treated as constant : \[\large z_x = ye^{2x^3-5} \cdot (6x^2) = 6x^2ye^{2x^3-5} \]

ganeshie8 (ganeshie8):

thats the partial with respect to x right ?

OpenStudy (anonymous):

\[z= xy^2, z _{x}= x2y \] i dont see why the y get multiplied by e^2x^3-5

OpenStudy (anonymous):

is it because of a property of e?

ganeshie8 (ganeshie8):

y can be treated as constant when u differentiate with respect to x, so u can simply pull it out of derivative

ganeshie8 (ganeshie8):

\[\large z = ye^{2x^3-5} \]

ganeshie8 (ganeshie8):

treat y as constant and differentiate with respect to x

ganeshie8 (ganeshie8):

\[\large \dfrac{\partial }{\partial x}(z) = \dfrac{\partial }{\partial x}\left[ye^{2x^3-5} \right]\]

ganeshie8 (ganeshie8):

\[\large \dfrac{\partial }{\partial x}(z) = y\dfrac{\partial }{\partial x}\left[e^{2x^3-5} \right]\]

ganeshie8 (ganeshie8):

\[\large \dfrac{\partial }{\partial x}(z) = y\left(e^{2x^3-5} \right) \dfrac{\partial }{\partial x}\left[2x^3-5\right] \]

ganeshie8 (ganeshie8):

\[\large \dfrac{\partial }{\partial x}(z) = y\left(e^{2x^3-5} \right) \left(6x^2\right)\]

ganeshie8 (ganeshie8):

\[\large \dfrac{\partial }{\partial x}(z) = 6x^2y\left(e^{2x^3-5} \right) \]

ganeshie8 (ganeshie8):

see if that makes more or less sense..